OpenStudy (anonymous):

the width of a rectangle is 7cm less than its length.if its perimeter is 50 cm,calculate its dimensions.

jhonyy9 (jhonyy9):

w=l-7 P=50 P=2l +2w =2l + 2(l-7) =2l +2l -14 = 4l-14 30=4l-14 44=4l l=11 w=11-7=4 w=4 l=11 P=22+8=30

jhonyy9 (jhonyy9):

sorry i have considered the perimeter 30

jhonyy9 (jhonyy9):

but the same you need to do for 50 too

jhonyy9 (jhonyy9):

50=4l-14 64=4l l=64/4 =16 l=16 w=16-7=9 l=16 w=9 P=32+18=50